OpenStudy (anonymous):

y=2x+1 find an equation for the inerverse relation

OpenStudy (anonymous):

this says multiply by 2, add 1. inverse is subtract 1, divide by 2 \[y=\frac{x-1}{2}\]

OpenStudy (anonymous):

or write \[x=2y+1\] and solve for y \[x=2y+1\] \[x-1=2y\] \[\frac{x-1}{2}=y\]